In a certain population of women 4 percent have had breast cancer (BC), 20 percent are smokers (S), and 3 percent are smokers and have had breast cancer (BC ∩S). A woman is selected at random from the population. What is the probability: P(BC ∪S)? [i.e. the probability that she has had breast cancer or smokes or both?
Answer #1

Using addition rule,

P(A or B) = P(A) + P(B) - P(A and B)

P(BC S) = P(BC) + P(S) - P(BC S)

= 0.04 + 0.20 - 0.03

= 0.21